OpenStudy (anonymous):

What is the d value in the following arithmetic sequence? (-7), (-2), 3, 8,... A. 5 B. 13 C. 3 D. (-2)

OpenStudy (anonymous):

well its increasing.

OpenStudy (anonymous):

-7 + x = -2 +7 +7 --- ---- x = +5

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

now check if +5 works...

OpenStudy (anonymous):

im just trying to double check my answers here.

OpenStudy (anonymous):

-7 + 5 = -2 -2 + 5 = 3 3 + 5 = 8 8 + 5 = 13 AND SO ON...

OpenStudy (anonymous):

well do u still think its B?

OpenStudy (anonymous):

@chicagoblkhaks ??

OpenStudy (anonymous):

i do. cause its increasing by 5 ...

OpenStudy (anonymous):

well if u still think its B, ur incorrect :) cuz be says 13... so idk what ur thinking lol

OpenStudy (anonymous):

*cuz B

OpenStudy (anonymous):

holy crap. its asking what its increasing by. not what the next number is. well im glad you caught that, cause im double checking all my questions in an assignments thats asking me a million different questions about athrimetic

OpenStudy (anonymous):

lol ;D np

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

good luck :)